OpenStudy (anonymous):

how would solve quadratic equations like x^2+2x-1=2 and 5r^2=80

OpenStudy (anonymous):

5r^2=80 r^2=80/5

OpenStudy (theviper):

To solve quadratic equation u should use the formula :- \[\LARGE{\color{green}{-b \pm \sqrt{b^2-4ac}\over2a}}\]

OpenStudy (anonymous):

so would r^2=80/5 be the answer or the first step

OpenStudy (theviper):

\[\LARGE{5r^2=80}\]\[\LARGE{r^2=\frac{80}{5}}\]\[\LARGE{r=\frac{\sqrt{80}}{\sqrt{5}}}\]

OpenStudy (callisto):

First: x^2+2x-1=2 x^2+2x-1-2=2-2 x^2+2x-3=0 Factor the expression: (x+3)(x-1) = 0 Put each factor equal to 0 and solve them. That is to solve: x+3 = 0 or x-1 = 0 Second one: Divide both sides by 5 first 5r^2 = 80 r^2 = 80/5 = 16 So, r^2 = 16 Take square root r= \(\pm \sqrt{16}\) = ...

OpenStudy (callisto):

If you want to use the quadratic formula mentioned in the above, make sure you've got the equation in the form \(ax^2 + bx + c = 0\)

OpenStudy (theviper):

yup right :)

OpenStudy (callisto):

@mopishmold2575 r^2 = 80/5 is only the first step.

OpenStudy (anonymous):

ok but how would you put it in the quadratic form since there 4 numbers youd have a number for the a variable the b variable and the c variable but what would you do with the fourth variable

OpenStudy (anonymous):

wait i got it thanks